OpenStudy (anonymous):

The area of Jared’s computer room floor is 168 ft. The formula for area of a rectangle is . If the length of the room is 14 ft., what is the width? a. 6 ft c. 12 ft b. 14 ft d. 17 ft

OpenStudy (anonymous):

If the room is a rectangle, it's just l*w. 14*x=168 x=168/14 x=12

OpenStudy (anonymous):

thanks you soooo much!!!

OpenStudy (anonymous):

No prob